If F=GMM/R^2. Find the dimension of G if R is the distance, F is the force and MM are the masses.
1
Expert's answer
2020-11-25T07:12:19-0500

Let's find the dimension of the universal gravitational constant, "G":


"G=\dfrac{FR^2}{M^2},""G=\dfrac{[M\dfrac{L}{T^2}][L]^2}{[M]^2}=M^{-1}L^3T^{-2}."

Answer:

"G=M^{-1}L^3T^{-2}."